OpenStudy (anonymous):

for the parallelogram, if m<2+5x-28 abd m<4=3x-10, find m< 3.

OpenStudy (anonymous):



OpenStudy (anonymous):

5x-28 = 3x-10 2x-28 = -10 2x = 18 x = 9. Substitute that into m<4... you get 3(9) - 10 = 17. This means angle 3 is...180-17 = 163.